A steel reinforcing rod is 15' long and 9/16" thick at its diameter. If a cubic foot of steel weighs 485 pounds, how much does the rod weigh?
Answer by checkley77(12844)

15*12(pi*9/32^2)=44.7 in^3 or 44.7/12*12*12=.0259ft^3*485=12.56lbs.